OpenStudy (anonymous):

what is the pattern? 1/80,1/40,1/20,1/10.

OpenStudy (anonymous):

More like multiply by 2.

OpenStudy (anonymous):

10*2=20 and 20*2=40 and 40*2=80 Multiply by 2

OpenStudy (anonymous):

\[\frac{ 1 }{ 80 }\times2=\frac{ 2 }{80 } or \frac{ 1 }{ 40}\] \[\frac{ 1 }{ 40 }\times2=\frac{ 2 }{ 40} or \frac{ 1 }{ 20}\]

OpenStudy (anonymous):

x2